OpenStudy (anonymous):

6x+9y=80 2x+20y=75 I need to solve for x & y using the elimination method.

OpenStudy (anonymous):

please show all steps

OpenStudy (anonymous):

6x+9y=80 (2x+20y=75)x3 6x+9y=80 6x+60y=225 subtract 6x+9y=80 -6x-60y=-225 = -51y=-145 y=(-145/-51) y=2.8431

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

what did you get for x?

OpenStudy (anonymous):

x = 925/102 and y = 145/51

OpenStudy (anonymous):

you put in y=2.8431 into one of the first equations. Here: when y=2.8431 6x + 9y = 80 6x = 80 - 9y x = (80 - 9y) / 6 x = (80 - 9*2.8431) /6 x = 9.06868